# Real-World Remote Sensing Image Dehazing: Benchmark and Baseline

### v1

@@ Abstract
Remote Sensing Image Dehazing (RSID) poses significant challenges in real-world scenarios due to the complex atmospheric conditions and severe color distortions that degrade image quality. The scarcity of real-world remote sensing hazy image pairs has compelled existing methods to rely primarily on synthetic datasets. However, these methods struggle with real-world applications due to the inherent domain gap between synthetic and real data. To address this, we introduce Real-World Remote Sensing Hazy Image Dataset (RRSHID), the first large-scale dataset featuring real-world hazy and dehazed image pairs across diverse atmospheric conditions. Based on this, we propose MCAF-Net, a novel framework tailored for real-world RSID. Its effectiveness arises from three innovative components: Multi-branch Feature Integration Block Aggregator (MFIBA), which enables robust feature extraction through cascaded integration blocks and parallel multi-branch processing; Color-Calibrated Self-Supervised Attention Module (CSAM), which mitigates complex color distortions via self-supervised learning and attention-guided refinement; and Multi-Scale Feature Adaptive Fusion Module (MFAFM), which integrates features effectively while preserving local details and global context. Extensive experiments validate that MCAF-Net demonstrates state-of-the-art performance in real-world RSID, while maintaining competitive performance on synthetic datasets. The introduction of RRSHID and MCAF-Net sets new benchmarks for real-world RSID research, advancing practical solutions for this complex task. The code and dataset are publicly available at here.
@@
Remote sensing, real-world dehazing, color calibration, multi-branch processing, adaptive fusion.
@@ I Introduction
Remote sensing (RS) imagery serves as a pivotal enabler for environmental sciences and agricultural applications, driving significant advancements in urban planning, agricultural monitoring, disaster management, and environmental conservation [1, 2, 3, 4, 5, 6]. However, the ubiquitous presence of atmospheric haze presents a persistent obstacle, degrading image quality by diminishing contrast and obscuring fine details. Such degradation severely compromises the precision and reliability of subsequent analytical tasks, underscoring the urgent need for efficient, robust dehazing algorithms tailored to practical RS applications.
@@
In recent years, deep learning has catalyzed remarkable progress in RS image dehazing (RSID) [8]. Pioneering efforts such as TransWeather [9] introduced Transformer architectures to dehazing, while 4KDehazing [10] advanced high-resolution dehazing through sophisticated multi-scale feature extraction. These methods have predominantly been developed and validated on synthetic datasets like RICE [11] and RS-HAZE [7], achieving impressive results—e.g., FFA-Net [12] with a PSNR of 33.52 dB on RICE, and DehazeFormer [7] with 39.57 dB on RS-HAZE. However, when deployed on real-world RS hazy images, these models frequently suffer substantial performance drops. This gap arises from three critical challenges:
@@
    Color Fidelity Challenges: Color Fidelity Challenges: Real-world hazy images display pronounced color distortions caused by multiple scattering and sensor-specific responses [15], as shown in Fig. 1, in stark contrast to the basic color shifts observed in synthetic datasets.
To address these challenges, we introduce the Real-World Remote Sensing Hazy Image Dataset (RRSHID). It is the first large-scale dataset featuring real-world hazy and dehazed image pairs across diverse atmospheric conditions and geographical contexts. Its construction is detailed in Section IV.
@@
Traditional RSID methods often adopt simplistic reconstruction techniques, such as direct feature mapping [17] or basic skip connections [18], neglecting the pivotal role of color calibration in real-world contexts. To address this, we introduce the Color-calibrated Self-supervised Attention Module (CSAM). CSAM combines self-supervised color learning with attention-driven refinement via two mechanisms: (1) an intermediate ”fake image” generated through a learnable color correction matrix for implicit color supervision, and (2) a query-key-value attention mechanism where color-calibrated features guide refinement through a global channel attention branch (for broad color consistency, e.g., sky regions) and a local depth-wise convolution branch (for detail preservation, e.g., urban edges). Tailored for RSID and inspired by recent restoration advances [19], CSAM bridges encoding and decoding, ensuring both color accuracy and structural integrity.
@@
    RRSHID: The first real-world RS hazy image dataset, encompassing real-world hazy and dehazed image pairs, fostering advancements in real-world RSID research.
@@
    MCAF-Net: A novel framework for real-world RSID, leveraging hierarchical feature refinement through specialized modules.
@@
    MFIBA: An efficient module for cross-dimensional feature integration via cascaded blocks and multi-branch processing, capturing haze-invariant patterns with computational efficiency.
@@
    CSAM: A color-calibration module addressing distortions and preserving details via self-supervised learning and attention-guided refinement.
  5. 5.
    MFAFM: An adaptive multi-scale fusion module preserving local details and global context for complex real-world scenarios.
@@ II-B Deep Learning-Based RSID
The advent of deep learning has transformed RSID, surpassing the limitations of traditional model-based techniques. Early contributions, such as DehazeNet [23] and MSCNN [24], pioneered convolutional neural networks (CNNs) for transmission map estimation, establishing a benchmark for superior performance. AOD-Net [25] advanced this paradigm by introducing end-to-end training for direct clean image recovery. DCPDN [26] further refined the approach by jointly optimizing transmission maps and atmospheric light within a cohesive framework. As the field matured, attention mechanisms and multi-scale architectures gained traction. FFA-Net [12] and GridDehazeNet [27] incorporated these features to effectively capture global context and local intricacies. PSMB-Net [28] extended this trajectory with dual-path feature interactions to tackle non-uniform haze in urban RS scenes, while MSBDN [29] introduced a boosted decoder for enhanced feature propagation. PSD-Net [30] employed pyramid supervision to enable progressive dehazing.
Recent innovations have shifted toward transformer-based architectures, marking a significant evolution in RSID. DehazeFormer [7] leveraged shifted window attention to model long-range dependencies, offering robust haze removal across complex scenes. PCSformer [31] built upon this by introducing cross-stripe attention modules to capture vertical and horizontal haze gradients, augmented by proxy tokens for computational efficiency. Trinity-Net [1] integrated gradient-guided Swin transformers with physics-informed priors, facilitating joint optimization of dehazing and downstream tasks, such as land cover classification.
To overcome data scarcity—a persistent challenge in RSID—hybrid learning strategies have emerged. SCANet [32] proposed a self-paced semi-curricular attention network, blending curriculum learning with channel-spatial attention to address non-homogeneous haze. Semi-supervised methods [33] harnessed both labeled and unlabeled data, while domain adaptation techniques [14] bridged synthetic and real-world domains. Self-supervised learning [34] has also shown promise in exploiting unlabeled RS data for improved dehazing. SFSNiD [35] explored frequency-consistent twin fusion, while MaxDehazeNet [36] applied maximum flow theory to model atmospheric scattering, enhancing physical interpretability.
@@ III Proposed Method
The encoder leverages MFIBA to perform efficient, cross-dimensional feature extraction through cascaded multi-branch processing, addressing the spatial heterogeneity of real-world haze. CSAM bridges the encoder and decoder by progressively refining color fidelity through attention-guided refinement, ensuring structural preservation. In the decoder, MFAFM adaptively fuses multi-scale features using dual attention mechanisms—channel attention for spectral recalibration and spatial attention for terrain-specific haze suppression—effectively balancing local details and global context. This design is tailored to tackle the complex atmospheric variability and color distortions prevalent in real-world RSID scenarios, outperforming conventional approaches reliant on static feature processing.
@@ III-A MFIBA
MFIBA processes input features $\mathbf{X}\in\mathbb{R}^{b\times c\times h\times w}$ by partitioning them into four subgroups $\{x_{1},x_{2},x_{3},x_{4}\}$, each with dimensions $(1,c/4,h,w)$. Three learnable queries—$\mathbf{Q}_{hw}$, $\mathbf{Q}_{ch}$ and $\mathbf{Q}_{cw}$—are dynamically generated for the spatial ($h$-$w$), channel-height ($c$-$h$), and channel-width ($c$-$w$) dimensions. These queries, initialized using Xavier normal distribution [39], interact with reshaped feature groups via Hadamard products $\mathbf{Q}_{ij}\odot\mathbf{X}_{ij}$, selectively attenuating redundant haze patterns while preserving cross-dimensional correlations. A cascaded architecture with three MFIB stages progressively refines contextual representations through residual connections, expanding the effective receptive field to capture large-scale haze structures. Computational efficiency is achieved through depth-wise separable convolutions [40] and 1$\times$1 channel interactions, reducing complexity from $\mathcal{O}(c^{2}hwk^{2})$ to $\mathcal{O}(c^{2}hw/4+3chw)$. This process is formalized as:
$\vspace{-1pt}\begin{split}&[X_{hw},X_{ch},X_{cw},X_{0}]=\text{LN}(\mathbf{X})\\ &\mathbf{Q}_{hw},\mathbf{Q}_{ch},\mathbf{Q}_{cw}=\text{DWConv}_{3\times 3}(\mathbf{Q})\\ &\mathbf{X}^{\prime}_{queried}=\text{Concat}(\mathbf{Q}_{ij}\odot\mathbf{X}_{ij}\text{Conv}_{1\times 1}(X_{0}))\\ &\mathbf{X}_{queried}=\text{MLP}(\mathbf{X}^{\prime}_{queried})+\mathbf{X}\end{split}\vspace{-1pt}$ (1)
@@ III-B CSAM
Conventional RSID [41, 42, 43] methods often overlook color calibration during early decoding stages [44], limiting their effectiveness against the severe color distortions [15] induced by atmospheric scattering in real-world scenarios. To address this, we introduce CSAM, integrated across multiple decoder levels, to progressively refine haze-free features while ensuring color accuracy and structural integrity, as shown in Fig. 4.
@@
By hierarchically correcting coarse-to-fine spectral distortions—global color balance in early stages (e.g., sky regions) and local chromatic consistency in later stages (e.g., vegetation)—CSAM effectively mitigates nonlinear atmospheric-sensor interactions, enhancing real-world RSID performance.
@@ III-C MFAFM
The resulting feature maps are then concatenated. To enhance channel-wise feature relationships, we employ a channel attention mechanism $\text{CA}(\cdot)$ Spatial relationships are further refined using a spatial attention mechanism $\text{SA}(\cdot)$ :
@@ IV RRSHID: A Real-World RSID Dataset
RSID is a persistent challenge, particularly in real-world contexts where atmospheric conditions exhibit significant complexity and variability. Although deep learning has driven notable progress in this domain, its application to authentic RS imagery remains constrained. Studies by Li et al. [37] and Qin et al. [12] demonstrate that existing methods often falter under diverse atmospheric scattering patterns and heterogeneous scene compositions. Furthermore, Zhang et al. [48] underscore the limited generalization of many models to real-world RS data, attributing this to the reliance on synthetic training datasets that fail to capture the intricacies of natural haze. To bridge this critical gap, we present the Real-World Remote Sensing Hazy Image Dataset (RRSHID), a pioneering collection of 3,053 paired hazy and haze-free images. Spanning urban, agricultural, and coastal landscapes across China, these images were captured under diverse illumination and seasonal conditions. Unlike synthetic datasets such as those in [11, 7, 49, 1], which depend on simplified atmospheric models, RRSHID leverages collaborative data acquisition with meteorological agencies. This approach ensures realistic representations of altitude-dependent haze distributions and sensor-induced spectral distortions, positioning RRSHID as an essential benchmark for evaluating dehazing algorithms and advancing domain generalization in real-world RSID.
@@ IV-A Image Acquisition and Processing
The development of the RRSHID involved a systematic methodology for image selection and preprocessing, implemented through collaboration with remote sensing domain experts. Working in concert with meteorological agencies and geospatial scientists, we employed multi-temporal satellite image alignment techniques to identify spatially congruent regions captured by satellite platforms across multiple acquisition cycles. This synergistic approach enabled the acquisition of radiometrically calibrated multi-temporal image pairs exhibiting identical geographic coordinates but distinct atmospheric profiles, thereby establishing a robust dataset for real-world RSID studies.
The dataset draws from the GF_PMS: L1A optical satellite, part of the Optical and SAR Satellite Payload Retrieval system, which delivers high-resolution imagery at 1-meter panchromatic and 4-meter multispectral scales. Covering diverse terrains across China—urban centers, agricultural fields, and coastal zones—the images were acquired 2021-2023, reflecting seasonal and illumination variability. Each raw satellite image, stored in GeoTIFF format with an approximate file size of 8 GB, underwent the following preprocessing pipeline:
1) Channel Conversion: Multispectral images with multiple bands were transformed into 3-channel RGB (red, green, blue) format using Adobe Photoshop, prioritizing bands optimal for visual analysis.
2) Image Cropping and Alignment: Overlapping regions in hazy and clear image pairs were identified and cropped with precision using QGIS (Quantum Geographic Information System), selected for its advanced geospatial capabilities and scalability with large satellite datasets.
3) Format Optimization: Custom Python scripts further refined the cropped images, generating $256\times 256$-pixel subimages and converting them from TIFF to PNG format to enhance compatibility with machine learning frameworks.

### v2

@@ Abstract
Remote Sensing Image Dehazing (RSID) poses significant challenges in real-world scenarios due to the complex atmospheric conditions and severe color distortions that degrade image quality. The scarcity of real-world remote sensing hazy image pairs has compelled existing methods to rely primarily on synthetic datasets. However, these methods struggle with real-world applications due to the inherent domain gap between synthetic and real data. To address this, we introduce Real-World Remote Sensing Hazy Image Dataset (RRSHID), the first large-scale dataset featuring real-world hazy and hazy-free image pairs across diverse atmospheric conditions. Based on this, we propose MCAF-Net, a novel framework tailored for real-world RSID. Its effectiveness arises from three innovative components: Multi-branch Feature Integration Block Aggregator (MFIBA), which enables robust feature extraction through cascaded integration blocks and parallel multi-branch processing; Color-Calibrated Self-Supervised Attention Module (CSAM), which mitigates complex color distortions via self-supervised learning and attention-guided refinement; and Multi-Scale Feature Adaptive Fusion Module (MFAFM), which integrates features effectively while preserving local details and global context. Extensive experiments validate that MCAF-Net demonstrates state-of-the-art performance in real-world RSID, while maintaining competitive performance on synthetic datasets. The introduction of RRSHID and MCAF-Net sets new benchmarks for real-world RSID research, advancing practical solutions for this complex task. The code and dataset are publicly available at https://github.com/lwCVer/RRSHID.
@@
Remote sensing, real-world dehazing, color distortion, multi-branch processing, adaptive fusion.
@@ I Introduction
Remote sensing (RS) imagery serves as a pivotal enabler for environmental sciences and agricultural applications, driving significant advancements in urban planning, agricultural monitoring, disaster management, and environmental conservation [1, 2, 3, 4, 5]. However, the ubiquitous presence of atmospheric haze presents a persistent obstacle, degrading image quality by diminishing contrast and obscuring fine details. Such degradation severely compromises the precision and reliability of subsequent analytical tasks, underscoring the urgent need for efficient, robust dehazing algorithms tailored to practical RS applications.
@@
In recent years, deep learning has catalyzed remarkable progress in RS image dehazing (RSID) [7]. Pioneering efforts such as TransWeather [8] introduced Transformer architectures to dehazing, while 4KDehazing [9] advanced high-resolution dehazing through sophisticated multi-scale feature extraction. These methods have predominantly been developed and validated on synthetic datasets like RICE [10] and RS-HAZE [6], achieving impressive results—e.g., FFA-Net [11] with a Peak Signal-to-Noise Ratio (PSNR) of 33.52 dB on RICE, and DehazeFormer [6] with 39.57 dB on RS-HAZE. However, when deployed on real-world RS hazy images, these models frequently suffer substantial performance drops. This gap arises from three critical challenges:
@@
    Color Distortion Challenges: Real-world hazy images display pronounced color distortions caused by multiple scattering and sensor-specific responses [14], as shown in Fig. 1, in stark contrast to the basic color shifts observed in synthetic datasets.
To address these challenges, we introduce the Real-World Remote Sensing Hazy Image Dataset (RRSHID). It is the first large-scale dataset featuring real-world hazy and hazy-free image pairs across diverse atmospheric conditions and geographical contexts. Its construction is detailed in Section IV.
@@
Traditional RSID methods often adopt simplistic reconstruction techniques, such as direct feature mapping [16] or basic skip connections [17], neglecting the pivotal role of color calibration in real-world contexts. To address this, we introduce the Color-calibrated Self-supervised Attention Module (CSAM). CSAM combines self-supervised color learning with attention-driven refinement via two mechanisms: (1) an intermediate ”fake image” generated through a learnable color correction matrix for implicit color supervision, and (2) a query-key-value attention mechanism where color-calibrated features guide refinement through a global channel attention branch (for broad color consistency, e.g., sky regions) and a local depth-wise convolution branch (for detail preservation, e.g., urban edges). Tailored for RSID and inspired by recent restoration advances [18], CSAM bridges encoder and decoder, ensuring both color accuracy and structural integrity.
@@
    RRSHID as the first real-world RS hazy image dataset, highlighting how it addresses fundamental limitations in existing synthetic datasets.
@@
    MCAF-Net specifically tailored to the challenges of real-world RSID, rather than simply adapting general image dehazing techniques.
@@
    Our proposed modules (MFIBA, CSAM, and MFAFM) that specifically address the unique characteristics of real-world RS scenarios, including complex haze distributions, color distortion issues, and the need to preserve both fine details and global context.
@@
    The practical significance of our approach for real-world RS applications, emphasizing computational efficiency and performance improvements in authentic scenarios.
@@ II-B Deep Learning-Based RSID
The advent of deep learning has transformed RSID, surpassing the limitations of traditional model-based techniques. Early contributions, such as DehazeNet [22] and MSCNN [23], pioneered convolutional neural networks (CNNs) for transmission map estimation, establishing a benchmark for superior performance. AOD-Net [24] advanced this paradigm by introducing end-to-end training for clean image recovery. As the field matured, attention mechanisms and multi-scale architectures gained traction. FFA-Net [11] and GridDehazeNet [25] incorporated these features to effectively capture global context and local intricacies. RSDehazeNet [26] and HyperDehazing [27] introduced attention mechanisms specifically designed for multispectral and hyperspectral RS images. PSMB-Net [28] extended this trajectory with dual-path feature interactions to tackle non-uniform haze in urban RS scenes, while DR3DF-Net [29] enhanced feature integration through dynamic routing, enabling adaptive fusion across scales.
Recent innovations have shifted toward transformer-based architectures, marking a significant evolution in RSID. DehazeFormer [6] leveraged shifted window attention to model long-range dependencies, offering robust haze removal across complex scenes. RSDformer [30] further enhanced image structure recovery by incorporating novel self-attention mechanisms to capture both local and global correlations. PCSformer [31] built upon this by introducing cross-stripe attention modules to capture vertical and horizontal haze gradients, augmented by proxy tokens for computational efficiency. Trinity-Net [1] integrated gradient-guided Swin transformers with physics-informed priors, facilitating joint optimization of dehazing and downstream tasks, such as land cover classification. Concurrently, state space models have emerged as compelling alternatives to traditional CNNs and transformer architectures. Notably, RSDehamba [32] and HDMba [33] pioneered the application of the Mamba architecture to RSID and hyperspectral RSID, respectively, demonstrating superior efficiency in modeling long-range dependencies with linear computational complexity.
To overcome data scarcity—a persistent challenge in RSID—hybrid learning strategies have emerged. SCANet [34] proposed a self-paced semi-curricular attention network, blending curriculum learning with channel-spatial attention to address non-homogeneous haze. Semi-supervised methods [35] harnessed both labeled and unlabeled data, while domain adaptation techniques [13] bridged synthetic and real-world domains. Self-supervised learning [36] has also shown promise in exploiting unlabeled RS data for improved dehazing. Recent unpaired training approaches have effectively addressed the synthetic-real domain gap. Diff-Dehazer [37] integrates diffusion priors with CycleGAN for bijective mapping, leveraging both physical priors and cross-modal representations. DNMGDT [38] utilizes a parameter-shared architecture with adaptive weighting of prior-based pseudo labels and physical model guided domain transfer to enhance generalization to real hazy scenes. SFSNiD [39] explored frequency-consistent twin fusion, while MaxDehazeNet [40] applied maximum flow theory to model atmospheric scattering, enhancing physical interpretability.
@@ III Proposed Method
The encoder leverages MFIBA to perform efficient, cross-dimensional feature extraction through cascaded multi-branch processing, addressing the spatial heterogeneity of real-world haze. CSAM bridges the encoder and decoder by progressively refining color distortion through attention-guided refinement, ensuring structural preservation. In the decoder, MFAFM adaptively fuses multi-scale features using dual attention mechanisms—channel attention for color channel recalibration and spatial attention for terrain-specific haze suppression—effectively balancing local details and global context. This design is tailored to tackle the complex atmospheric variability and color distortions prevalent in real-world RSID scenarios, outperforming conventional approaches reliant on static feature processing.
@@ III-A MFIBA
MFIBA processes input features $\mathbf{X}\in\mathbb{R}^{b\times c\times h\times w}$ by partitioning them into four subgroups $\{x_{1},x_{2},x_{3},x_{4}\}$, each with dimensions $(1,c/4,h,w)$. Three learnable queries—$\mathbf{Q}_{hw}$, $\mathbf{Q}_{ch}$ and $\mathbf{Q}_{cw}$—are dynamically generated for the spatial ($h$-$w$), channel-height ($c$-$h$), and channel-width ($c$-$w$) dimensions. All queries are independently randomly initialized. In multi-branch partial queried learning, we transform ${x_{1}}$, ${x_{2}}$, and ${x_{3}}$ into formats $(b,c,h,w)$, $(b,w,c,h)$, and $(b,h,c,w)$, respectively. We then apply bilinear interpolation to adjust the last two dimensions of queries to match feature groups, and employ depth-wise convolutions to enhance query representations. These queries interact with reshaped feature groups via Hadamard products $\mathbf{Q}_{ij}\odot\mathbf{X}_{ij}$, selectively attenuating redundant haze patterns while preserving cross-dimensional correlations. For ${x_{4}}$, we solely apply 1$\times$1 convolution for channel-wise information interaction, achieving partial self-relevant feature representation. Subsequently, the four groups of enhanced features are concatenated and transformed through a simple feed-forward network. A cascaded architecture with three MFIBs stages progressively refines contextual representations through residual connections, expanding the effective receptive field to capture large-scale haze structures. Computational efficiency is achieved through depth-wise separable convolutions [43] and 1$\times$1 channel interactions, reducing complexity from $\mathcal{O}(c^{2}hwk^{2})$ to $\mathcal{O}(c^{2}hw/4+3chw)$. This process is formalized as:
$\vspace{-1pt}\begin{split}&[X_{hw},X_{ch},X_{cw},X_{0}]=\text{LN}(\mathbf{X})\\ &\mathbf{Q}_{hw},\mathbf{Q}_{ch},\mathbf{Q}_{cw}=\text{DWConv}_{3\times 3}(\mathbf{Q})\\ &\mathbf{X}^{\prime}_{queried}=\text{Concat}(\mathbf{Q}_{ij}\odot\mathbf{X}_{ij},\text{Conv}_{1\times 1}(X_{0}))\\ &\mathbf{X}_{queried}=\text{MLP}(\mathbf{X}^{\prime}_{queried})+\mathbf{X}\end{split}\vspace{-1pt}$ (1)
@@ III-B CSAM
RS images under hazy conditions often exhibit severe color distortion, which is more complex compared to near-ground images, as atmospheric thickness and composition vary significantly at different altitudes. Conventional RSID [44, 45, 46] methods often overlook color calibration during early decoding stages [47], limiting their effectiveness against the severe color distortions [14] induced by atmospheric scattering in real-world scenarios. To address this, we introduce CSAM, integrated across multiple decoder levels, to progressively refine haze-free features while ensuring color accuracy and structural integrity, as shown in Fig. 4.
@@
Our proposed CSAM module offers two significant contributions. First, it operates specifically in the color space with a learnable correction matrix that adaptively adjusts color distribution based on input characteristics, addressing complex color distortion in haze-affected RS imagery. Second, CSAM implements an innovative self-attention mechanism that differs from similar strategies optimized for object tracking[50, 51] by: (1) focusing on color rather than spatial features to address differential atmospheric scattering effects across color channels; (2) maintaining multi-scale color consistency through MFIBA module integration; and (3) specifically targeting real-world RS dehazing challenges rather than general vision tasks. These design elements enable CSAM to effectively address color restoration challenges in RSID, achieving superior color correction while preserving critical image details.
@@ III-C MFAFM
The resulting feature maps are then concatenated. To enhance channel-wise feature relationships, we employ a channel attention mechanism $\text{CA}(\cdot)$ that recalibrates RGB channel relationships to address color distortion. Spatial relationships, which identifies and processes regions with varying haze densities across different terrain types, are further refined using a spatial attention mechanism $\text{SA}(\cdot)$:
@@ IV RRSHID: A Real-World RSID Dataset
RSID is a persistent challenge, particularly in real-world contexts where atmospheric conditions exhibit significant complexity and variability. While deep learning has driven notable progress in this domain, its application to authentic RS imagery remains constrained. Studies by Li et al. [41] and Qin et al. [11] demonstrate that existing methods often falter under diverse atmospheric scattering patterns and heterogeneous scene compositions. Furthermore, Zhang et al. [53] underscore the limited generalization of many models to real-world RS data, attributing this to the reliance on synthetic training datasets that fail to capture the intricacies of natural haze. To bridge this critical gap, we present the Real-World Remote Sensing Hazy Image Dataset (RRSHID), a pioneering collection of 3,053 paired hazy and haze-free images. spanning urban, agricultural, and coastal landscapes across China, including provinces such as Anhui, Sichuan, Shanghai, Shandong, and Fujian, these images were captured under diverse illumination and seasonal conditions. Unlike synthetic datasets such as those in [10, 6, 54, 1], which depend on simplified atmospheric models, RRSHID leverages collaborative data acquisition with meteorological agencies. RRSHID ensures realistic representations of altitude-dependent haze distributions and sensor-induced color distortions. It serves as a vital benchmark for evaluating dehazing algorithms and advancing domain generalization in real-world RSID.
@@ IV-A Image Acquisition and Processing
The development of RRSHID involved a systematic methodology for image selection and preprocessing, implemented through collaboration with remote sensing domain experts. Working in concert with meteorological agencies and geospatial scientists, we employed multi-temporal satellite image alignment techniques to identify spatially congruent regions captured by satellite platforms across multiple acquisition cycles. To ensure the authenticity and precision of our paired images, we developed a comprehensive dataset generation framework consisting of three key stages:
IV-A1 Data Source Selection
The dataset draws from the GF_PMS: L1A optical satellite, part of the Optical and SAR Satellite Payload Retrieval system, which delivers high-resolution imagery at 1-meter panchromatic and 4-meter multispectral scales. Covering diverse terrains across China—urban centers, agricultural fields, and coastal zones—the images were acquired 2021-2023, reflecting seasonal and illumination variability. Each raw satellite image, stored in GeoTIFF format with an approximate file size of 8 GB, contains rich metadata including acquisition time, sensor parameters, and precise geolocation information.
IV-A2 Temporal-Spatial Alignment
The critical challenge in creating paired hazy/haze-free images lies in ensuring precise alignment between degraded and clear images. Our approach addresses this by a dual temporal-spatial alignment strategy:
  • Temporal Proximity Selection: We first identified temporally proximate multispectral image pairs (typically within 1-3 months) from the same geographic region that exhibited contrasting atmospheric conditions—one with significant haze presence and one with clear visibility. This temporal proximity minimizes seasonal variations in vegetation, building structures, and other land features.
  • Geospatial Coordinate Matching: For each identified image pair, we implemented precise geospatial alignment by extracting geographic coordinate metadata (latitude/longitude) from image headers and identifying overlapping regions with consistent land features but different atmospheric conditions. Through precise marking of the top-left and bottom-right geographic coordinates of target regions, we ensured exact spatial correspondence before applying coordinate-based cropping to extract identical geographic areas from both images.
IV-A3 Image Preprocessing
Following alignment, each image pair underwent a standardized preprocessing workflow:
  • Channel Conversion: Multispectral images with multiple bands were transformed into 3-channel RGB (red, green, blue) format using Adobe Photoshop, prioritizing bands optimal for visual analysis.
  • Format Optimization: Custom Python scripts refined the cropped images, generating $256\times 256$ pixels subimages and converting them from TIFF to PNG format to enhance compatibility with machine learning frameworks.
TABLE I: Comparison of RRSHID with Existing Remote Sensing Datasets
Dataset Size Real/Synthetic Paired Resolution Application Color Bias Image Type Scene Diversity
StateHaze1K [54] 1,200 Synthetic Yes 512$\times$512 Dehazing No RGB Single
RSID [1] 1,000 Synthetic Yes 256$\times$256 Dehazing No RGB Single
RS-HAZE [6] 51,300 Synthetic Yes 512$\times$512 Dehazing No RGB Single
LHID [55] 31,017 Synthetic Yes 512$\times$512 Dehazing No RGB Multiple
DHID [55] 14,990 Synthetic Yes 512$\times$512 Dehazing No RGB Single
HyperDehazing [27] 2,000 Synthetic Yes 512$\times$512 Dehazing No Hyperspectral Multiple
HyperDehazing [27] 70 Real No 512$\times$512 Dehazing No Hyperspectral Multiple
RICE1 [10] 500 Real Yes 512$\times$512 Dehazing No RGB Single
RICE2 [10] 736 Real Yes 512$\times$512 Cloud/Haze Removal No RGB Single
HRSI [56] 796 Real No 512$\times$512 to 4000$\times$4000 Dehazing (Object Detection) No RGB Single
RRSHID 3,053 Real Yes 256$\times$256 Dehazing Yes RGB Multiple
Figure 6: Haze density distribution across the RRSHID dataset.